40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog, Page 40 of 56  Not logged in ELOG logo
ID Date Author Type Category Subject
  838   Tue Jun 29 00:40:53 2010 DmassLaserGYROMode Matching

When posting MATLAB code to the elog - attach as an archive so it doesn't take up many pages when viewing all posts in the elog

  837   Mon Jun 28 17:33:32 2010 JennaLaserGYROMode Matching

 Here is the code I've been working on for mode matching to the cavity. MM.m is adapted from goo2.m from Rana's page (http://www.ligo.caltech.edu/~rana/mat/MM/PMCmatch/), and there you can also find the functions called by MM.m that I didn't edit, such as chaindo. It finds the optimum focal lengths and distances for the lenses.

EtaCalcSimple calculates the overlap function for the beam assuming a non-astigmatic beam and an astigmatic cavity. 

MMplot plots the values for z (distance from the waist) and w as a function of the changing distance from the mirror to the first lens.

Calculate allows you to change values such as the distances to see how this affects the values for eta, the waist size, and z.

***EDIT  Alastair Heptonstall ***

I've taken all of the .m files and put them in an archive file.

Attachment 1: mode_match.tar
  836   Mon Jun 28 16:13:21 2010 AidanComputingGeneralcurrently no access to ATF from outside

 

Done.


Quote:

the network connection is down again, so our router has to be restarted ...

whoever will go to the lab on Monday first plz powercycle the linksys router...

 

  835   Mon Jun 28 01:36:57 2010 FrankComputingGeneralcurrently no access to ATF from outside

the network connection is down again, so our router has to be restarted ...

whoever will go to the lab on Monday first plz powercycle the linksys router...

  834   Sun Jun 27 18:02:51 2010 DmassComputingGeneralOven interfacing

Quote:

Here is the in loop temperature for various loop settings.

Legend: The first number is proportional gain, the F is integral gain ("fast" as called by the 3040) To note:

  • I don't know why the loop oscillates for lower gain settings
  • The loop seems most stable for the highest possible gain settings ("300 Fast"), so I will use this.
  • This is actually Kelvin/rtHz

 

 Making sense of this - I spent a couple minutes sitting down and writing down the right equation for the complete open loop transfer function. Assuming an integral gain of "fast", we have:

OLTF = 3.68e-3*Kp*(s+8.5/Kp)/s/(s+0.0167)

Where Kp is the proportional gain. When we make Kp 50, we put the zero from the PI loop at 0.17, where the pole is at 0.0167. This makes a nice long chunk of f^-2 in the loop, murdering the phase. As we crank the gain up to 300, we push this zero down, and shorten the stretch of f^-2.

Ideally with an integral gain of "fast" we would have a proportional gain of ~500.

If we go to "slow", then we have for our complete open loop transfer function:

OLTF = 3.68e-3*Kp*(s+1.6/Kp)/s/(s+0.0167)

For a gain of "1 slow", as in the plot, we find the zero is 2 orders higher than the pole, which is consistent with this setting ringing way more than the others. A gain of "100 slow" would work for balancing the zero and pole. Our highest proportional gain available with a "slow" setting is 50, which is similar to the situation above, where the zero is about a factor of 2 before the pole, thought with a lower UGF than above.

 

  833   Thu Jun 24 21:18:49 2010 DmassComputingGeneralOven interfacing

Here is the in loop temperature for various loop settings.

Legend: The first number is proportional gain, the F is integral gain ("fast" as called by the 3040) To note:

  • I don't know why the loop oscillates for lower gain settings
  • The loop seems most stable for the highest possible gain settings ("300 Fast"), so I will use this.
  • This is actually Kelvin/rtHz

 

Attachment 1: OvenLoopTesting.pdf
OvenLoopTesting.pdf
  832   Thu Jun 24 18:49:04 2010 FrankMiscGeneralUSB-to-serial converter back at 40m

The USB-to-serial converter is back at 40m.
The  new ones Rana bought are not working properly so the got the one from the ATF which we use for the particle counter.

  831   Wed Jun 23 21:53:50 2010 DmassComputingGeneralOven interfacing

 

File info for temperature
Filename Proportional Gain Integral Gain
OLtemp off off
CLtemp1 50 fast
CLtemp2 100 fast
CLtemp3 300 fast
CLtemp4 50 (longer time) fast

 

  830   Wed Jun 23 18:15:59 2010 AidanLab InfrastructureHVACNew temperature controllers

We now have two temperature controllers in the lab:

A couple of weeks ago they installed a second temperature controller on the South Wall. This drives the HVAC heater that is above the stereo.

The original temperature controller (West Wall) was also upgraded to have a fancy new 1960's-style mechanical display of the setpoint-needle on the front type. Maybe in 50 or 60 years we can get a digital controller in here.

At the moment the second controller has not been calibrated to match the original controller.

  1. 00001.jpg - new controller on South Wall
  2. 00002.jpg - close up of South Wall controller
  3. 00003.jpg - upgraded West Wall controller
Attachment 1: 00001.jpg
00001.jpg
Attachment 2: 00002.jpg
00002.jpg
Attachment 3: 00003.jpg
00003.jpg
  829   Wed Jun 23 17:57:00 2010 FrankLab InfrastructureHVACtemperature fluctuations - before and after

the two graphs show the temperature fluctuation in the ATF for 14 days each. The first graph from March, the second for the last 14 days.
scale is the same for both graphs.
Couldn't plot a single graph for the last couple of month as the dataviewer stops displaying most of the June data if you want to plot more than 40 consecutive days or so even if it tells you to do it right.

 

temp_ATF1.png

temp_ATF2.png

  828   Wed Jun 23 15:59:13 2010 DmassComputingGeneralOven interfacing

This is a measurement of the open loop temperature of the oven - the blue box is "off", so this may be higher than the temperature noise the servo normally has to supress.

  • It is worth noting that the readout of the box has 0.0076 K precision (this is the difference between values).
  • I don't know if it does this (12/14 bit?) rounding in the servo loop or not

THIS IS POWER!! - that is sad. take a square root of the spectrum to translate it into amplitude spectral density - using this as a clipboard because I somehow don't have pwelch on my Ubuntu MATLAB install o.O. That is also sad.

ovenCL50a=load('CLtemp1.txt');
ovenCL100=load('CLtemp2.txt');
ovenCL300=load('CLtemp3.txt');
ovenCL50b=load('CLtemp4.txt');
ovenOL=load('OLtemp.txt');

Fs=1;
nfft=200*Fs;
[pwrOL,frOL]=pwelch(detrend(ovenOL),hanning(nfft),nfft/2,nfft,Fs);
Fs=4; % sample frequency in Hz
nfft=200*Fs;
[pwr50b,fr50b]=pwelch(detrend(ovenCL50b),hanning(nfft),nfft/2,nfft,Fs);
[pwr100,fr100]=pwelch(detrend(ovenCL100),hanning(nfft),nfft/2,nfft,Fs);
[pwr300,fr300]=pwelch(detrend(ovenCL300),hanning(nfft),nfft/2,nfft,Fs);


figure(1)
loglog(...
    frOL,sqrt(pwrOL),'r',...
    fr50b,sqrt(pwr50b),'k',...
    fr100,sqrt(pwr100),'cyan',...
    fr300,sqrt(pwr300),'magenta',...
    'LineWidth',1)
axis tight
grid

Attachment 1: OvenOLMeas.pdf
OvenOLMeas.pdf
  827   Wed Jun 23 13:47:01 2010 FrankLab InfrastructureHVAClatest particle counts
  • the dates where they tried to fix the leaks are around/after the first peak in particle counts (04/23/2010)
  • i don't have data for a few days after that because nobody saved the data from the counter i guess and it starts to overwrite it after some while (the white gap).
  • the latest peak from last week was due to some technical issues with the HVAC system which they fixed immediately (06/17/2010)

AdhikariLab_data_061810.png

  826   Tue Jun 22 21:53:24 2010 DmassComputingGeneralOven interfacing

edited the code to the following. This will save the data ~ 1x / second (I don't really care about a slight cumalative error in this since it's constant - it'll just be a systematic error in my frequency scaling)

 

global SERPORT
set SERPORT [open "/dev/ttyS0" "w+"]

fconfigure $SERPORT -mode 9600,n,8,1 -handshake none -buffering line

proc run_query { command } {
global SERPORT

puts $SERPORT "$command"
flush $SERPORT
gets $SERPORT line

return $line
}

#@puts [run_query "BEEP?"]

         set filename "/users/dmass/scripts/test.txt"
         set fileId [open $filename "w"]

 for { set i 1 } { $i <= 2000 } { incr i } {
#       puts "[clock format [clock seconds]]\t[run_query "TEC:T?"]"
                puts $fileId [run_query "TEC:T?"]
        after 1000
        }
close $fileId

  825   Tue Jun 22 19:17:27 2010 DmassComputingGeneralOven interfacing

I asked Vladimir in passing how he preferred to talk to things via the serial port, and he wrote this. It is now contained in /users/dmass/scripts/serial_test.tcl :

 

#!/usr/bin/env tclsh

global SERPORT
set SERPORT [open "/dev/ttyS0" "w+"]

fconfigure $SERPORT -mode 9600,n,8,1 -handshake none -buffering line

proc run_query { command } {
global SERPORT

puts $SERPORT "$command"
flush $SERPORT
gets $SERPORT line

return $line
}

#@puts [run_query "BEEP?"]

while { 1 } {

    puts "[clock format [clock seconds]]\t[run_query "TEC:T?"]"
    after 100
    }

  824   Mon Jun 21 20:46:40 2010 DmassComputingGeneralOven interfacing

To talk to the Temperature controller via RS 232, use

> sudo minicom

set this via the minicom software

  • baud rate = 9600
  • no parity
  • 8 data bits
  • 1 stop bit
  • Then do
  • E for echo on
  • do "BEEP?", should get 0 or 1
  • Reset the modem if you don't get a response
  • do "BEEP?" again, success, we are talking to the newport 3040

Next up - have a script dump the response to a temperature query every second or so to a file. This is inferior to using some AD590s to sense the temperature via the front end, because of the time constants involved in loading / reading out files. I will buy some AD590s for future use (though they will get here after I am finished with this).

Useful command(s)

  • TEC:T? - queries the temperature of the TEC sensor (and RTD in my case)
  • TEC:SET:T? - queries the temperature set point
  • TEC:T xxxxxx - can set the temperature setpoint
  • sudo stty -F /dev/ttyS0 -parenb echo

I was having problems because /dev/ttyS0 was set to root permissions, and a sudo wouldn't cut it when I tried to echo something to it. I changed the permissions and this now works:

  • echo "TEC:T?">/dev/ttyS0
  823   Mon Jun 21 09:44:31 2010 FrankElectronicsGYROIFR / Marconi Phase noise w/ Rubidium clock lock

very nice measurement!
how large was the external modulation input span? disabled?
if disabled, did you measure the phase noise with the input enabled (but terminated) and set to a usefull range as well?

RA: c.f. the entry below for html link to description.

  822   Sat Jun 19 19:20:29 2010 ranaElectronicsGYROIFR / Marconi Phase noise w/ Rubidium clock lock

There is a 10x improvement in the Marconi phase noise spectrum when locking it to the FS725 Rubidium clock.

To see if we could improve on the Marconi spectra for use in the Gyro/RefCav measurements I set up two of the 40m Marconis for "Direct External 10 MHz Reference" lock.

In the attached plot you see the BLUE trace (free running Marconis) is higher than the PURPLE (locked to Rubidium) trace. The BROWN dots are the SSB phase noise specs

for the FS725 Rubidium clock. In this case, I have the FS725 locked to the 1PPS from our GPS receiver, but I think it doesn't have any effect above 10 mHz.


Some Notes:

  1. The 'Direct' lock mode of the IFR2023A implies that it locks directly to the external reference. The 'Indirect' mode means that it first locks an internal TCXO to the external reference and then uses the TCXO as the internal reference. Since the FS725 has a better SSB phase noise spec than the IFR2023's internal PLL, I used the 'Direct'.
  2. As you can tell from the noise improvement, it looks like the internal PLL has a ~1kHz UGF but only has a gain of ~10 below there. That's pretty strange if true. We should purchase another Rubidium clock to beat with this one to see if its really as good as its spec.
  3. As usual, the beat spectra shown are the quadrature sum of 2 oscillators, so divide by sqrt(2) to estimate the noise of an individual IFR/Marconi.
  4. In all my previous phase noise measurements, I had neglected to compensate for the 30 mHz AC coupling of the SR560 I am using as a pre-ADC preamp. I've corrected for this in the calibration of these measurements. Its almost a negligible effect but I include it anyway for my internal satisfaction.
  5. Because of the high input impedance of the 10 MHz inputs of the IFRs, I have daisy-chained the 2 by using a BNC T on the back of the first one. The cable connecting to the bottom one is terminated with an in-line 50 Ohm terminator.
  6. As described previously in the 40m elog, the measurement setup is using a ZP-3MH mixer as a phase detector. Looking at the peak-peak signal with free running oscillators allows us to calibrate the phase detector slope when locked. An AC coupled SR560 with G=1000 is used as the preamp between the mixer and the ADC. Between the mixer's IF output and the SR560, there is a 50 Ohm inline term and a 1.9 MHz BNC low pass filter from Mini-Circuits.
  7. Since the FS725 spec is better than the PURPLE trace, I am assuming that the PURPLE measurement is real. If the FS725 was noisier, we would be fooling ourselves by cancelling out the common-mode noise of the FS725, but in our case we are dominated by the internal noise of the Marconis.
  8. Simply taking the FS725's 10 MHz output and deriving a 95 MHz or 160 MHz signal wouldn't gain us much, since the phase noise would also increase by the multiplication factor. For the Refcav experiment, the best option may be to bring the beat down into a reasonable regime by using 2 lasers rather than an AOM. Then we may be able to put the beat frequency below the Nyquist of the digital system and forego any RF demodulation.
Attachment 1: IFR.png
IFR.png
  821   Fri Jun 18 18:07:53 2010 DmassComputingDAQAttempt at oven transfer functions

This was wrong.

  820   Fri Jun 18 18:04:23 2010 DmassComputingDAQAttempt at oven transfer functions

Quote:

And a measurement of the oven's step response

  • I used a lab power supply as a current source for the TEC on the oven
  • I tweaked the current nob to make a current step
  • I looked at the step response
  • I used a phone video of the readout on the front of the Newport 3040 to get time stamps b/c it was so easy
  • (I will still be hooking up the RS232)

Results

  • I get a time constant of 60 seconds
  • This is an oven pole frequency of 16.7 mHz
  • The gain is 0.368

 

 

  819   Fri Jun 18 00:48:11 2010 ranaLab InfrastructureHVACHVAC now working again

I bought about a million converters - Koji has most of them, but I left one with Jan so that you can take that one to get the particle data.  Along with the HVAC, this will be a combination of low temperature and particle physics. Very advanced.

  818   Thu Jun 17 15:14:58 2010 DmassComputingDAQAttempt at oven transfer functions

And a measurement of the oven's step response

  • I used a lab power supply as a current source for the TEC on the oven
  • I tweaked the current nob to make a current step
  • I looked at the step response
  • I used a phone video of the readout on the front of the Newport 3040 to get time stamps b/c it was so easy
  • (I will still be hooking up the RS232)

Results

  • I get a time constant of 60 seconds
  • This is an oven pole frequency of 16.7 mHz

 

Attachment 1: CuOvenIStepResponse.pdf
CuOvenIStepResponse.pdf
  817   Thu Jun 17 11:05:42 2010 FrankLab InfrastructureHVACHVAC now working again

a minute ago a guy showed up in the lab and wanted to talk to Aidan but couldn't find him. He told me that the air conditioning is now working again and i should let him know.

Particle count in the ATF was about 1.5million before, now decreasing and already below 100k. Can't post data as the usb-to-serial converter is still used at 40m...

  816   Wed Jun 16 23:10:28 2010 DmassComputingDAQAttempt at oven transfer functions

Quote:

3040 P/I settings -

I = Slow <=> 0.0162 A/s/K

P = 2 <=> 0.0172 A/K

 

 A Couple more measurements yielded:

  • I = Slow <=> 0.0172 A/s/K
  • P = 1 <=> 0.0098 A/K

And

  • I = Fast <=>0.0848 A/s/K

Next I am hooking up the RS232 connector to one of the computers to see if I can get temperature logged, so that I can actually do something with it.

  815   Wed Jun 16 15:54:20 2010 DmassComputingDAQAttempt at oven transfer functions

3040 P/I settings -

I = Slow <=> 0.0162 A/s/K

P = 2 <=> 0.0172 A/K

 

  814   Tue Jun 15 23:47:09 2010 DmassElectronicsDoublingOven Control Updates

After talking to Koji, and learning a little about how digital integrators worked, I understood the wording of the manual differently. I now believe it to mean

"I have a true integrator with two gain settings, and a proportional term with ~12"

  813   Mon Jun 14 14:40:10 2010 FrankMiscGeneralcompany for temp ctrls and diode drivers

http://www.teamwavelength.com

e.g. these things is not entirely useless

http://www.teamwavelength.com/products/product.asp?part=19

http://www.teamwavelength.com/products/product.asp?part=58

http://www.teamwavelength.com/products/product.asp?part=29

  812   Fri Jun 11 21:27:50 2010 KojiComputingGeneralX11 on WS1 can not find nVidia driver

Dmass and Koji

We rebooted ws1 during the work, and started not to have X11 working.
So far the single monitor (with cloning) is working.

The machine claimed that it could not find the driver named nvidia.
I switched the driver to the generic driver (VESA Generic Driver) using a diagnostic dialog
which appeard when we could not run the X11 during the boot.

Using the generic driver we now can display 1280x1024 but can not use the dual head function.
I removed the existing nvidia driver (which also did not work) in order to reinstall it.
Then I tried to reinstall it, but the machine could not download it from the fedra site.

I will investigate the driver situation for nVidia GeForce 7600 GT and will try to install the alternative driver later.

 

  811   Fri Jun 11 20:33:17 2010 DmassComputingDAQAttempt at oven transfer functions

The plot I attached and didn't explain is:

  • Green = Voltage across the Ohm Ranger as driven and read out by the (unknown) bridge of the Newport 3040
  • Red = Voltage across a high power 5 Ohm resistor I am using on the current output of the Newport 3040
  • Units are standard DAQ conversion ~ 1600 counts / Volt

Notes:

  • I use an SR560 to buffer the signal from the ohm ranger before inputting it into the AA chassis
  • The resistance between the inner BNC leads of the AA box was ~2 kOhms - FOR SEVERAL DIFFERENT CHANNELS
  • I have the PI (no D) servo set at its lowest settings, so lowest integral gain, and lowest proportional gain.
  • I make step responses in the green curve via the ohm ranger.
  • I should see step responses of the red from the proportional term, and changes in the slope of the integration following these step responses.
  • I see SLOPE IN THE GREEN. This is weird.
  • The slope in the green goes away when I unplug the voltage (from the TEC drive current) across the 5 ohm resistor from the front end.
  • The slope in the green stops when the current rails.
  • I don't understand the coupling, but since I don't have circuit diagrams for the 3040 - I stopped thinking about it, and just used a USB stick in a scope to take this data.

Moral:

  • The 3040/ front end acts funny when you try to read out the current and the voltage of the thermometer.
  • I don't really care that this is the case.
  • The front end seems to be working close to all the way
Attachment 1: dmassFE.png
dmassFE.png
  810   Fri Jun 11 02:44:09 2010 FrankComputingDAQFB0 fixed

i think i have a loose connection on the timing pcb in the blue chassis in the psl lab. I had this "no sync" problem the day before i left too and after i checked all connections and wiggled all the cables everything was fine when i checked again. So try to do that with the 1Hz BNC sync cable connected at the back of the chassis. The psl model uses only the first card in the chassis and was working. I will replace the board beginning next week. If one of the models doesn't sync the framebuilder doesn't start.

Another option: Because we need both systems running in order to work with it i will try to use the bad frontend computer at a lower sampling rate. The spikes in the timing are about 280us. If the model is running at 2k we have about 480us cycle time, so it shouldn't matter anymore. I've mounted everything in the rack already to try that but had no time to look into details...

Quote:

The front end (fb0) was freezing at "Starting udev" every time I tried to boot it. I googled a tad and found some people somewhere said that this happened to them when some cables were loose, so I wiggled all the cables and rebooted.

FB0 now boots.

I think something may have been jiggled loose from the back of fb0 when the rack was manhandled, possibly during the fixing of the ceiling leaks.

  • I sshed into fb0
  • I did "startatf" and "startpsl"
  • I set both burtrestores to 1 -
  • Nothing worked yet - framebuilder kept respawing at 5 minute intervals
  • I checked the log files and noticed nothing obvious
  • I added a channel in the C2ATF.ini because I wasn't sure there was one
  • I restarted both processes again
  • The ATF model up with both ADC and DAC channels started working
  • I am unsure if the PSL model is working, the GDS_TP screen says "NO SYNC" under sync source.

 

  809   Thu Jun 10 21:18:18 2010 DmassComputingDAQFB0 fixed

The front end (fb0) was freezing at "Starting udev" every time I tried to boot it. I googled a tad and found some people somewhere said that this happened to them when some cables were loose, so I wiggled all the cables and rebooted.

FB0 now boots.

I think something may have been jiggled loose from the back of fb0 when the rack was manhandled, possibly during the fixing of the ceiling leaks.

  • I sshed into fb0
  • I did "startatf" and "startpsl"
  • I set both burtrestores to 1 -
  • Nothing worked yet - framebuilder kept respawing at 5 minute intervals
  • I checked the log files and noticed nothing obvious
  • I added a channel in the C2ATF.ini because I wasn't sure there was one
  • I restarted both processes again
  • The ATF model up with both ADC and DAC channels started working
  • I am unsure if the PSL model is working, the GDS_TP screen says "NO SYNC" under sync source.
  808   Thu Jun 10 02:08:15 2010 FrankLab InfrastructureGeneralseveral things offline / not working as usual

thanks

Quote:

PEM power cycled at around 11am this morning.

Quote:

several devices seem to be offline since the rack has been moved., e.g. i can't connect to the PEM device mounted on the wall above the rack (the little white box with the antenna), no connection to the OPC server is possible (the 1HU computer) and fb0 is broken

 

can someone power cycle the PEM device (white box) for me plz. I will take care of the rest when i'm back...

and can someone move the particle counter next to WS1 or WS2 and connect it to one of the serial ports using the special cable which we use to download the collected data (it's labeled) (plz don't forget to move the power supply too) 

thanks

 

 

  807   Wed Jun 9 15:13:15 2010 AlastairLab InfrastructureGeneralseveral things offline / not working as usual

PEM power cycled at around 11am this morning.

Quote:

several devices seem to be offline since the rack has been moved., e.g. i can't connect to the PEM device mounted on the wall above the rack (the little white box with the antenna), no connection to the OPC server is possible (the 1HU computer) and fb0 is broken

 

can someone power cycle the PEM device (white box) for me plz. I will take care of the rest when i'm back...

and can someone move the particle counter next to WS1 or WS2 and connect it to one of the serial ports using the special cable which we use to download the collected data (it's labeled) (plz don't forget to move the power supply too) 

thanks

 

  806   Tue Jun 8 10:49:34 2010 FrankLab InfrastructureGeneralseveral things offline / not working as usual

several devices seem to be offline since the rack has been moved., e.g. i can't connect to the PEM device mounted on the wall above the rack (the little white box with the antenna), no connection to the OPC server is possible (the 1HU computer) and fb0 is broken

 

can someone power cycle the PEM device (white box) for me plz. I will take care of the rest when i'm back...

and can someone move the particle counter next to WS1 or WS2 and connect it to one of the serial ports using the special cable which we use to download the collected data (it's labeled) (plz don't forget to move the power supply too) 

thanks

  805   Tue Jun 8 10:35:16 2010 AidanMiscstuff happensMoved some equipment to 40m

I've taken the following items to the 40m.

  1. 40MHz AOM
  2. Thorlabs Broadband EOM
  3. 2x New Focus 9071 4-axis stages
  4. ZHL-1A RF amplifier
  5. ZHL-2A RF amplifier

They're all marked "Adhikari Lab".

  804   Fri Jun 4 23:22:41 2010 DmassElectronicsDoublingOven Control Updates

I have turned to the SHG ovens:

  • I noticed a lot of extra non twisted pair of homemade cabling between my ovens and the Newport 3040 so I redid the cables to both the ovens.
  • (N.B. the Newport 3040 seems to have been discontinued, or at least is no longer searchable through google or the newport site)
  • The manual said to use a 15 pin D sub with the following connections (twisted pairs on +/- pairs):
Pin

Purpose

My Wiring
1 TEC+ N/C
2 TEC+ To TEC
3 TEC- To TEC
4 TEC- N/C
5 Ground Ground @ back of 3040
6 Ground Ground @ back of 3040
7 Sensor + To RTD
8 Sensor - To RTD
13 Interlock Send N/C
15 Interlock return N/C
  • I hooked up each cable to the respective ovens, and followed the 3040 manuals instructions to do a "RTD Null" calibration - this is just calibrating out wire resistance for some display number, I am unsure if this does anything to the servo
  • I noticed one of the temperature controllers did not servo the two displayed temperatures together
    • Under Menu > System > Calibration > Set T Null ; doing the forward arrow twice changed this to null displayed temperature
  • Both ovens now seem to be servoing the temperature to the setpoint.

Next - Figure out how to set these crappy temperature controllers up so that they are less crappy-

  • It seems I have about 10 Proportional gain settings, but....
  • I have NO Intergral gain setting, just a "slow" and "fast" setting, which seems to be the low frequency pole of the integrator according to the manual.
  • Yes, that's correct, it seems like I can only move the integral pole between two frequencies, and am not allowed to change the gain.
  • I shall explore this 20 item parameter space later.

 

  803   Fri Jun 4 21:32:54 2010 FrankComputingDAQfb0 reboot

Power cycled again via fron power button - couldn't get a command line up to do a proper reboot. Again frozen at "Starting udev"

 

I will leave it like this on the off chance that it is just taking a long time.

  802   Fri Jun 4 13:52:44 2010 FrankComputingDAQfb0 reboot

Quote:

can someone reboot fb0 plz.

 POower cycled once, had no command line.

 

Stuck on "starting udev" - do I keep power cycling this thing?

  801   Fri Jun 4 13:46:26 2010 DmassElectronicsDoublingThe Box Itself

  Edited Box Picture in post this one replies to

  800   Thu Jun 3 18:29:10 2010 Alastair & RanaLaserGYROMode matching and cavity alignment

 Today Rana bolted down the cavity mirrors and did some alignment.  I measured the distances between mirrors and have worked out a rough mode matching scheme.  I'll post the beamscan measurements separately since they will also be useful for the final mode matching calculation for the astigmatism in the cavity (and beam).  We probably want to make a beam measurement for the second path as well after it comes out of the AOM double-pass loop.

I've put the lenses in place for the rough mode matching and started aligning the cavity again.  The astigmatism in the beam is clearly worse than it was before because of the 6m radius curved mirrors.

Edit --- I've spent a while trying to get the mode aligned.  It still looks pretty messy though there is a 00 trying to break through every now and again.  I'm not sure why it looks so screwy - the light around the edges just seems to be in random patterns which we didn't see before.  I wonder if this is just light being lost because of the astigmatism?

  799   Thu Jun 3 15:49:06 2010 FrankComputingDAQfb0 still not working, plz reboot it - can't work since a week

[controls@fb1 ~]$ ssh fb0
ssh_exchange_identification: Connection closed by remote host
[controls@fb1 ~]$

  798   Thu Jun 3 10:59:37 2010 FrankComputingDAQfb0 reboot

can someone reboot fb0 plz.

  797   Thu Jun 3 03:24:41 2010 DmassLab InfrastructureFuglyRack terrible-ness

The rack has been moved away from the wall, and the 300V power supply to the PZT is half perched on/in the blue ADC box, whose top is partially hanging into the box itself. This was very much not the state it was last seen in.

Attachment 1: 00001.jpg
00001.jpg
  796   Thu Jun 3 03:23:09 2010 DmassComputingDAQfb0 down?

Quote:

plz, can someone check what's going on with fb0. I can't login but ping is working. workstations and fb1 are ok

 

I turned the monitor switch to fb0 on the rack, and saw nothing come up on screen. I presume this means it needs a hard reset?

DO WE HAVE ANY IDEA  WHY FB0 SEEMS TO NEED A HARD RESET AT A RATE OF ABOUT 1 / MONTH? THIS SEEMS TERRIBLE. MAYBE WE ARE BUMPING SOME RACK CONNECTION FREQUENTLY? Is this why we are burning through hard drives like toast?

  795   Wed Jun 2 20:24:26 2010 rana, alastairMiscGYROOptics moved

Today we moved some of the optics away from the blue box to make room for the mode matching.

We also found a 1" diameter optic in one of the cabinets with a transmission of 0.81 % (S-pol). This makes a nice match with the 0.57% transmission of the 2" diameter output coupler and so its now installed. We'll later replace it with a good 2" diameter mirror.

We also moved all of the optics after the laser to align with the screw holes and are trying to make the beams go along the screw hole lines.

Alastair has remeasured the laser beam profile after the EOM and the PBSs using the WinCamD beam scanner. Results are being produced.

  794   Wed Jun 2 17:59:57 2010 ranaLab InfrastructureGeneralparticle counter + lab temp

 

the manual how to use the particle counter is where the rest of the manuals are. If you want to use it read the manual BEFORE you touch it and reconfigure everything so that nothing is working anymore. If you still don't know how to use it or can't find the manual you can ask me or simply google for it. If you only want to read the latest values you don't have to program or even touch the thing at all. Simply connect it to a computer via serial port. Right now it is taking data every 10min or so for a short period of time. Up to 8000 measurements are stored in the device, starting to overwrite the oldest ones when reaching the limit.

The software is installed on the small laptop, the same we use for the WinCam. There is an icon on the desktop, i think the software is called "comet" or so. You have to connect the cable with the USB-to-serial adapter. The cable has a label "particle counter". Plug everything together,  run the software, and from the menu simply choose something like read everything, meaning data and settings. That's it.

Data is saved in comma-separated  textfile, using current date and time. Sometimes it happens that it can't plot the data later on because the data is corrupt. I don't know why this happens but it simply forgets some characters while reading from the serial port or saves it twice. As it is a simple textfile this is easy to fix.

 The Particle Counter is a GT-526 from Met One. Datasheet on the web here.

The manual is here.

I was unable to get the data off of it using my Mac and Terminal. I tried to get the data off of it using Frank's wonderfully helpful elog entry, but was still unable to. When retrieving data using the 'Comet' software on the little laptop, the window just says '0% complete' and continues counting. After 12 minutes, I gave up.

 

The setup of the counter seems to be that it adds up for 10 seconds and then waits for 590 seconds. The display on the front panel is in units of 'CF".

The current reading (while Alastair and I are working on the table) is:

0.3 micron     30000 CF

0.5 micron      3000 CF

1.0 micron      1700 CF

I'm assuming that we multiply by 6 to get to the commonly used CFM (# of particles per cubic foot per minute).


Around 5:30PM today, Alastair commented on how the 2 HVAC systems seemed to be fighting. Recently, the physical plant guys installed a second AC sensor (stay tuned for the elog and photos from Aidan). Apparently, these systems don't feed the same controller. So I decreased the temperature on the new South wall sensor by 1.5 F down to 68 F. The one in the NW corner of the lab is unchanged.

  793   Wed Jun 2 17:11:58 2010 ZachLaserGYROMirror transmission

Ah.. right. I remember that this possibility dawned on me when I went upstairs to calculate finesses that day. I meant to tell you that I didn't check that the polarization was still S---sorry.

Quote:

I re-measured the mirror transmissions for 45S for all the candidate mirrors for the cavity.  The Y1 45P mirrors that were borrowed from the 40m have very low transmission for 45S.  I think that when they were measured before, the polarization had been rotated at the wave plate just after the faraday, and this is what gave the higher transmission values for the 45P mirrors.  I rotated it back to S using one of the good polarizers and power meter.  I have labelled each mirror with a number and the measured transmission.

 

Mirror 1:  BS1 - 99 - 45S, 2".  Old input mirror from the gyro.  Power in 115.4mW.  Power transmitted 0.657mW.  Transmission 0.57%

Mirror 2: Y1 - 45S, 2".  Old output mirror from the gyro.  Power in 115.3mW.  Power transmitted 19uW.  Transmission 165ppm.

Mirror 3: Y1 - 45P, 2".  From the 40m.  Power in 115.1mW.  Power transmitted 13uW.  Transmission 113ppm.

MIrror 4 Y1 - 45P, 2".  From the 40m.  Power in 115.3mW.  Power transmitted 20.2uW.  Transmission 175ppm.

Mirror 5: Y1 - 45S - 6.00m, 1".  Old curved mirror from gyro.  Power in 114.9mW.  Power transmitted 15.3uW.  Transmission 133ppm.

Mirror 6: Y1 - 45S - 6.00m, 1".  Old curved mirror from gyro.  Power in 115mW.  Power transmitted 20.5uW.  Transmission 178ppm.

 

  792   Wed Jun 2 14:05:59 2010 AlastairLaserGYROMirror transmission

I tried some more mirrors:

Mirror 7 : CVI PR1-1064-99.5-1037 for 0degrees, 1".  Transmission 1.28% for 45S.  Taken from the PSL lab.  Mirror is dirty.

Mirror 8 : CVI PR1-1064-98-1037 for 0degrees, 1".  Transmission 69% for 45S.  Taken from PSL lab.  Mirror is dirty.

Mirror 9 : VLOC FM-25.4-6.35-1FS-1010-1064-2, 1".  Transmission 101ppm.  Taken from 40m.

One possible configuration would be to use Mirror 7 as the input and Mirror 1 as the output.  That gives us a finesse of around 330 and a contrast defect of 0.13.


Quote:

I re-measured the mirror transmissions for 45S for all the candidate mirrors for the cavity.  The Y1 45P mirrors that were borrowed from the 40m have very low transmission for 45S.  I think that when they were measured before, the polarization had been rotated at the wave plate just after the faraday, and this is what gave the higher transmission values for the 45P mirrors.  I rotated it back to S using one of the good polarizers and power meter.  I have labelled each mirror with a number and the measured transmission.

 

Mirror 1:  BS1 - 99 - 45S, 2".  Old input mirror from the gyro.  Power in 115.4mW.  Power transmitted 0.657mW.  Transmission 0.57%

Mirror 2: Y1 - 45S, 2".  Old output mirror from the gyro.  Power in 115.3mW.  Power transmitted 19uW.  Transmission 165ppm.

Mirror 3: Y1 - 45P, 2".  From the 40m.  Power in 115.1mW.  Power transmitted 13uW.  Transmission 113ppm.

MIrror 4 Y1 - 45P, 2".  From the 40m.  Power in 115.3mW.  Power transmitted 20.2uW.  Transmission 175ppm.

Mirror 5: Y1 - 45S - 6.00m, 1".  Old curved mirror from gyro.  Power in 114.9mW.  Power transmitted 15.3uW.  Transmission 133ppm.

Mirror 6: Y1 - 45S - 6.00m, 1".  Old curved mirror from gyro.  Power in 115mW.  Power transmitted 20.5uW.  Transmission 178ppm.

 

  791   Wed Jun 2 11:20:02 2010 AlastairLaserGYROMirror transmission

I re-measured the mirror transmissions for 45S for all the candidate mirrors for the cavity.  The Y1 45P mirrors that were borrowed from the 40m have very low transmission for 45S.  I think that when they were measured before, the polarization had been rotated at the wave plate just after the faraday, and this is what gave the higher transmission values for the 45P mirrors.  I rotated it back to S using one of the good polarizers and power meter.  I have labelled each mirror with a number and the measured transmission.

 

Mirror 1:  BS1 - 99 - 45S, 2".  Old input mirror from the gyro.  Power in 115.4mW.  Power transmitted 0.657mW.  Transmission 0.57%

Mirror 2: Y1 - 45S, 2".  Old output mirror from the gyro.  Power in 115.3mW.  Power transmitted 19uW.  Transmission 165ppm.

Mirror 3: Y1 - 45P, 2".  From the 40m.  Power in 115.1mW.  Power transmitted 13uW.  Transmission 113ppm.

MIrror 4 Y1 - 45P, 2".  From the 40m.  Power in 115.3mW.  Power transmitted 20.2uW.  Transmission 175ppm.

Mirror 5: Y1 - 45S - 6.00m, 1".  Old curved mirror from gyro.  Power in 114.9mW.  Power transmitted 15.3uW.  Transmission 133ppm.

Mirror 6: Y1 - 45S - 6.00m, 1".  Old curved mirror from gyro.  Power in 115mW.  Power transmitted 20.5uW.  Transmission 178ppm.

  790   Wed Jun 2 09:37:03 2010 AlastairLab InfrastructureGeneralParticle count

~80,000

  789   Wed Jun 2 07:16:38 2010 FrankComputingDAQfb0 down?

plz, can someone check what's going on with fb0. I can't login but ping is working. workstations and fb1 are ok

ELOG V3.1.3-